Please help me solve this problem.
Two angles are supplementary. The measure of one is twice the measure of the other. What is the measure of the larger angle?
Thanks.

Answer by checkley71(8403)

X+Y=180
X=2Y NOW REPLACE X BY 2Y & SOLVE FOR Y
2Y+Y=180
3Y=180
Y=180/3
Y=60 ANSWER FOR THE SMALLER ANGLE.
X=2*60=120 ANSWER FOR THE LARGER ANGLE.
PROOF:
60=120=180
180=180